OpenStudy (anonymous):

What is the function rule for the following situation? Rex paid $20 for a membership to the pool and pays $3.00 each time he goes to the pool. a. y=20x+3 b. 20=3x+y c. y=x+20 d. none of these

OpenStudy (shelby1290):

D. because the correct answer is y = 20 + 3x

OpenStudy (shelby1290):

The reason being is this- look at y as the total amount of money Rex pays to use the pool. He paid a one time fee of $20 which covers his membership, and an additional $3 for every time he visits. So look at x as his number of visits, the correct equation becomes y = 20 + 3x. For example he pays the $20 to become a member, then he visits the pool 3 times the next week. The total amount he has now paid becomes $20 + $3*3, so y comes out to be 20 + 9 = $29.